Fabulously fit offers a membership for $35 per month plus a $50 enrollment fee. The fitness studio offers memberships for $40 per month plus a $35 enrollemnt fee. in how many months will the fitness clubs cost the same? what will the cost be?
Answer by mananth(16949)

let x = number months
Fabulously fit= 50+35x
Fitness studio 35+40x
50+35x=35+40x
5x=15
x= 3 months